OpenStudy (anonymous):

What is $1800 dollars at 2% for ten years?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (wolf1728):

formula is TOTAL = Principal * (1+rate)^n where n is number of years TOTAL = 1,800.00 * (1.02)^10 TOTAL = 1,800.00 * 1.21899442 TOTAL = 2,194.19 Here's a calculator to check your work: http://www.1728.org/compint.htm